OpenStudy (anonymous):

find the locus of P(x,y) such that the sum of the square of its distance from (-1,0) and (1,0)

OpenStudy (helder_edwin):

the sentence is imcoplete.

OpenStudy (helder_edwin):

*incomplete

OpenStudy (anonymous):

is 26 sorry

OpenStudy (anonymous):

|(x-1)^2+y^2|+|(x+1)^2+y^2\=26?

OpenStudy (helder_edwin):

let A(-1,0) and B(1,0) and P(x,y). if i understand it is \[ \large [d(P,A)+d(P,B)]^2=26 \] am i wrong?

OpenStudy (anonymous):

isnt it ((d(P,A))^2 + other one?

OpenStudy (helder_edwin):

i don't know. the wording of the problem is confusing

OpenStudy (anonymous):

sum so + of square of dist idkk

OpenStudy (anonymous):

yes... it should be: \(\large d(P,A)^2+d(P,B)^2=26 \)

OpenStudy (anonymous):

so just absolute valcue?

OpenStudy (anonymous):

hyelpp??

OpenStudy (helder_edwin):

well if @dpaInc is correct, then using the definition of distance \[ \large 26=d(P,A)^2+d(P,B)^2= \] \[ \large =[(x+1)^2+(y-0)^2]+[(x-1)^2+(y-0)^2] \]

OpenStudy (anonymous):

okay so just do it normally??

OpenStudy (helder_edwin):

yes